# DeStein: Navigating Detoxification of Language Models via Universal Steering Pairs and Head-wise Activation Fusion

### v1

@@ 4.1 Experimental setting
Datasets. For our experiments, we use the RealToxicityPrompts (RTP) dataset(Gehman et al., 2020), comprising 100K text segments. Each segment’s beginning serves as a prompt, with toxicity scores annotated using the Perspective API. To ensure fairness, we re-evaluate these scores as Pozzobon et al. (2023a) suggests. See Appendix A for more details. According to updated scoring, we sort 5k prompts with toxicity scores below 0.5 for non-toxic prompts and the rest for toxic prompts. Due to significant computational demands, we randomly select 1k prompts from each category for experiments with LLMs.
Models. In our research, we employ the GPT2-large(Radford et al., 2019a) as our foundational model to establish a baseline for our experiments. However, to demonstrate the scalability of our approach across LLMs, we also present results utilizing different model families, including LLaMA-2-7b(Touvron et al., 2023), OPT-6.7b(Zhang et al., 2022), and MPT-7b(Team, 2023).
@@
Baselines. For the GPT2-large, we compare our approach against two fine-tuning methodologies: DAPT and DisCup, in addition to three decoding-based methods, namely GeDi, DEXPERTS, and GOODTRIEVER. Furthermore, we explore a prompt-based method, Self-Detoxify. For LLMs, our baseline comparison includes Self-Detoxify and a decoding-based method named LMA. More details can be seen in Appendix B.
@@
Implementation Details. To ensure a fair comparison, our experimental setup for generation follows the practices established by Liu et al. (2021), employing the most commonly used nucleus sampling to generate 25 continuations for each prompt. For GPT2-large, we use $\alpha_{contr}$=0.38 and $l$=20. For LLMs, we use $\alpha_{contr}$=0.3 and $l$=40. More details are described in Appendix D.
4.2 Automatic evaluation
@@
In our research, we evaluate generated text on three key aspects: toxicity, fluency, and diversity, following established methodologies(Gehman et al., 2020). (1) Toxicity: We utilize the toxicity score from the Perspective API, which encompasses both Expected Maximum Toxicity (Exp. Max. Tox.) and Toxicity Probability (Tox. Prob.). (2) Fluency: we compute perplexity (PPL) using a slightly larger model from the same family. For instance, the PPL of GPT2-large’s generations are calculated using GPT2-XL. (3) Diversity: we calculate the mean of distance $n$-grams.
@@
Results on GPT2. The results of the automatic evaluation, as shown in Table 1, indicate that DeStein significantly outperforms all existing baselines in detoxification. Notably, DeStein is among the most fluent methods while also preserving output diversity to a great extent compared to the base model. Additionally, our method does not require any gradient information from the model during the inference process and solely operates on activations, thus not introducing significant inference time overhead. The generation runtime for each method is reported in Table 2 to demonstrate the efficiency of DeStein relative to other methods. Additionally, we provide further analysis of toxic and non-toxic prompts in Appendix F.
Table 1: Automatic evaluation results on RTP dataset with GPT2-large. Best models are shown in bold and 2nd best models are underlined.
Category Model Toxicity ($\downarrow$) Fluency ($\downarrow$) Diversity ($\uparrow$)
Exp. Max. Tox. Tox. Prob. PPL Dist-1 Dist-2 Dist-3
Base Model GPT2-large 0.557 0.567 27.252 0.588 0.856 0.850
@@
Decoding-based GeDi 0.416 0.314 67.595 0.579 0.856 0.852
@@
Prompt-based Self-Detoxify 0.360 0.235 40.689 0.584 0.868 0.862
OURs 0.203 0.061 37.809 0.574 0.860 0.860
Table 2: Inference time involves producing 20-token continuations for each of the 100 prompts on a 4090 GPU.
Model Inference Time (s) ($\downarrow$) Relative to GPT2 (large) ($\downarrow$) Parameter Count
GPT2-large 6.134 774M
OURs 7.013 +14% 774M
Self-Detoxify 10.583 +73% 774M
Dexperts 21.237 +246% 3 $\times$ 774M
Results on LLMs. Moreover, to demonstrate the effectiveness and scalability of our method in LLMs, we conducted detoxification experiments across various families of LLMs, with the automatic evaluation results presented in Table 3.
Table 3: Automatic evaluation results on RTP dataset with LLMs. Best models are shown in bold.
Model Toxicity ($\downarrow$) Fluency ($\downarrow$) Diversity ($\uparrow$)
Exp. Max. Tox. Tox. Prob. PPL Dist-1 Dist-2 Dist-3
LLaMA-2-7b 0.539 0.550 17.687 0.612 0.851 0.828
Self-Detoxify 0.413 0.318 83.972 0.648 0.876 0.839
@@
OURs 0.296 0.170 29.160 0.618 0.858 0.835
OPT-6.7b 0.622 0.661 16.127 0.565 0.839 0.841
Self-Detoxify 0.559 0.554 75.019 0.582 0.864 0.856
@@
OURs 0.291 0.157 17.733 0.562 0.850 0.855
Applying the Self-Detoxify method to LLMs, we observe a significant reduction in detoxification effectiveness and a noticeable decrease in fluency. The LMA method, designed specifically for LLMs, does not achieve satisfactory detoxification results. Due to the different decoding methods employed by LMA compared to the basic decoding method in this paper, it is not fair to directly compare fluency and diversity. This can be attributed to the immense parameter size of LLMs, with the Self-Detoxify method, oriented towards smaller-scale models (e.g., GPT2-large), failing to adapt to LLMs, and the classifier-dependent LMA method also failing to achieve efficient detoxification. However, without fine-tuning or auxiliary models, our method still demonstrates competitive detoxification effects across these categories of LLMs, and our method performs well across different model families, proving its effectiveness and scalability.
4.3 Black-box LLMs evaluation
@@
Relying exclusively on automated metrics for evaluation falls short of being thoroughly comprehensive. Therefore, we incorporate black-box LLMs as supplementary evaluators, specifically including GPT-4 and Gemini(Durmus et al., 2022). We compare our method’s generative outputs against top-performing baselines: DisCup and DEXPERTS for GPT2-large, and Self-Detoxify and LMA for LLMs. For design details about the prompt, please see Appendix C. Based on our findings that the number of prompts did not significantly affect the results, we extracted a balanced test dataset with 50 toxic and 50 non-toxic prompts for evaluation. The evaluation results indicate that our method exhibits lower toxicity and improved fluency compared to the best-performing baselines. The results are presented in Table 4.
@@
Table 4: Black-box LLMs Evaluation Results. Win, Lose, and Draw respectively represent the percentage of times our method outperforms, underperforms, or equals the baseline.
Model Versus GPT-4 Gemini
Win Lose Draw Win Lose Draw
GPT2-large DisCup 0.72 0.28 0.00 0.64 0.21 0.15
DEXPERTS 0.79 0.21 0.00 0.63 0.21 0.16
LLaMA-2-7b Self-Detoxify 0.71 0.29 0.00 0.63 0.23 0.14
LMA 0.74 0.25 0.01 0.64 0.19 0.17
5 Analysis
@@ 5.1 Analysis on the influence of detoxification strength
Figure 2: Trade-off between detoxification strength and PPL.
For detoxification methods during the inference stage, when the control intensity exceeds a certain critical value, it often leads to a rapid decrease in the fluency of generated text, rendering the text completely unusable. This limitation hinders the effectiveness of many methods in achieving efficient detoxification. However, the approach proposed in this paper detoxifies layer by layer through representations in the activation space, combined with probe techniques to apply different control intensities at different positions, which can largely avoid this problem. Our experiments, with results in Figure 2, demonstrate this. We take the difference between the toxicity scores of the base model and the detoxified model using our method as the detoxification strength and observe the relationship between fluency and detoxification strength. After reaching point A, the model’s generative capabilities only begin to deteriorate significantly. However, by the time the control intensity reaches point A, the toxicity has already been reduced to 0.030. Such a low level of toxicity becomes imperceptible to humans. The range of control intensity before point A is already sufficient to meet the model’s detoxification needs, so the generative collapse beyond point A does not have an impact.
Previous decoding methods were largely based on the logit space, and as the control intensity increased, generation collapse became inevitable. The success of our experiment provides a new perspective for controllable generation that more efficient control may be achievable in the activation space. This aligns with the intuition behind controllable generation methods based on VAEs(Sohn et al., 2015).
5.2 Analysis on interpretability in the activation space
Controllable generation methods based on VAE leverage a low-dimensional bottleneck to learn a disentangled latent space, while LMs often struggle to achieve good controllability at the activation layer due to the complexity of high-dimensional spaces. This proposed method mitigates the complexity of high-dimensional spaces to some extent through steering pairs and head-wise fusion. In this section, we verify the interpretability of the method by analyzing the data distribution in the activation space.
@@ 5.2 Analysis on interpretability in the activation space
As illustrated in Figure 3(a), we employed heatmap to visualize the accuracy of linear classifiers for each attention head across all layers, to analyze the effects and underlying principles of Probing-based Weights.
To further bolster the credibility of our method, we investigate the existence of a toxicity-non-toxicity vector within the activation space. Specifically, we focus on the activation space corresponding to the 6th attention head of the 23rd layer, as illustrated in Figure 3(b). This layer and head were selected due to their high classification accuracy identified through heatmap analysis in Probes. Employing the t-SNE(Maaten & Hinton, 2008), we visualize the dimensionality-reduced representation of toxic and non-toxic sentences within this activation space. Our results reaffirm the presence of a linear representation between toxicity and non-toxicity in the activation space. Moreover, to more vividly demonstrate the distribution differences in various activation spaces, we also visualize the space associated with the 12th layer and 7th head, as illustrated in Figure 3(c), which exhibits a classification accuracy nearly equivalent to that of random selection in the heatmap analysis. The distribution of toxic and non-toxic sentences within these two spaces further validates our method’s theoretical soundness. Additionally, we conducted comprehensive ablation experiments on attention heads, with details provided in Appendix E.
@@ 6 Conclusions
In this paper, we propose a novel approach to toxicity mitigation, named DeStein, which is based on the activation space for detoxification. Specifically, our method utilizes the self-induced data of the model to identify the detoxification vectors. During the inference phase, the detoxification vectors and the output of the module are merged, and probe techniques are employed to apply weights to head-wise activations. Experimental results demonstrate that our approach not only achieves efficient detoxification but also preserves the generative capacity of the model to the greatest extent possible. Notably, DeStein maintains considerable performance in toxicity mitigation without significantly increasing inference time. Unlike previous methods, DeStein exhibits flexibility and competitiveness when dealing with LLMs and is scalable across different model families.
@@ 7 Limitaions
Our methodology, predicated on the assumption of linear representation, employs arithmetic operations to decouple toxic attributes and general capabilities, yet this approach is notably constrained. Initially, constructing an ideal parallel pair poses considerable difficulty. Moreover, this method represents an indirect form of decoupling, facing theoretical challenges in achieving complete separation. Consequently, there is a necessity to explore more efficient and versatile decoupling strategies, such as those based on causal reasoning, knowledge guidance, or meta-learning techniques. These methods aim to effectively unify general capabilities with safety, delineating our subsequent direction of research.

### v3

@@ 4.1 Experimental settings
Datasets. We use the RealToxicityPrompts (RTP) dataset (Gehman et al., 2020), comprising 100K text segments. Each segment’s beginning serves as a prompt, with toxicity scores annotated using the Perspective API. For fairness, we re-evaluate these scores as Pozzobon et al. (2023a) suggests. See Appendix A for more details. According to the updated scores, we classify prompts with scores below 0.5 as non-toxic and the rest as toxic. We randomly selected 5k prompts per category for experiments with GPT2-large, and 1k prompts for LLMs.
Models. Following the previous studies, we applied detoxification techniques to GPT2-large (Radford et al., 2019a). Moreover, to demonstrate the scalability of our approach across a range of LLMs, we expanded our analysis to include various model families: GPT2-XL (1.3B), LLaMA2 (7B and 13B) (Touvron et al., 2023), OPT (6.7B) (Zhang et al., 2022), and MPT (7B) (Team, 2023).
@@
Baselines. For GPT2-large, we compare our approach against two finetuning-based methods: DAPT and DisCup, and four finetuning-free methods, namely GeDi, DEXPERTS, GOODTRIEVER and Self-Detoxify. For LLMs, our baseline comparison includes Self-Detoxify and a decoding-based method called LMA. More details are provided in Appendix B.
@@
Implementation Details. For fair comparison, our experimental setup for generation follows the practices established by Liu et al. (2021), employing the most commonly used nucleus sampling to generate 25 continuations for each prompt. For GPT2-large, we use $\alpha_{contr}$=0.38 and $m$=20. For LLMs, we use $\alpha_{contr}$=0.3 and $m$=40. Further details are described in Appendix D.
4.2 Evaluation results
We employed both statistical metrics and the LLM-as-a-Judge approach to enhance the persuasiveness of our results.
@@ 4.2 Evaluation results
Main results. For statistical metrics, we evaluate generated text on three key aspects: toxicity, fluency, and diversity, following established methods (Gehman et al., 2020). (1) Toxicity: We utilize the toxicity score from the Perspective API, which encompasses both Expected Maximum Toxicity (EMT) and Toxicity Probability (TP). (2) Fluency: we compute perplexity (PPL) using a slightly larger model from the same family. (3) Diversity: we calculate the mean of distance $n$-grams.
@@
The results, as shown in Table 1, indicate that DeStein significantly outperforms existing methods. Notably, DeStein is among the most fluent methods while also preserving output diversity to a great extent compared to the base model. Additionally, our method does not require any gradient information from the model during the inference process and solely operates on activations, thus not introducing significant inference time overhead. The generation runtime for each method is reported in Table 2 to demonstrate the efficiency of DeStein relative to other methods.
Type Method Toxicity $\downarrow$ Fluency $\downarrow$ Diversity $\uparrow$
EMT TP PPL Dist-1 Dist-2 Dist-3
- Base 0.557 0.567 27.252 0.588 0.856 0.850
@@
Finetuning-free GeDi 0.416 0.314 67.595 0.579 0.856 0.852
@@
Self-Detoxify 0.360 0.235 40.689 0.584 0.868 0.862
DeStein 0.203 0.061 37.809 0.574 0.860 0.860
Table 1: Evaluation results on the RTP dataset with GPT2-large. The best results are shown in bold, and the 2nd best results are underlined. The finetuning-free method is divided into two parts by a dividing line: the upper part requires support from auxiliary models or databases, while the lower part does not.
Method Inference Time $\downarrow$ Time Increase $\downarrow$ Parameter
Base 6.134s 774M
DeStein 7.013s +14% 774M+$\epsilon$
Self-Detoxify 10.583s +73% 774M
DEXPERTS 21.237s +246% 3 $\times$ 774M
Table 2: Inference time involves producing 20-token continuations for each of the 100 prompts on a 4090 GPU with GPT2-large. $\epsilon$ is a small positive constant, negligible in comparison to the model’s parameter count. Detailed calculation is in Appendix E.
Moreover, to demonstrate the effectiveness and scalability of our method in LLMs, we conducted detoxification experiments across various families of LLMs, with the automatic evaluation results presented in Table 3. Applying the Self-Detoxify method to LLMs, we observe a significant reduction in detoxification effectiveness and a noticeable decrease in fluency. The LMA method, designed specifically for LLMs, does not achieve satisfactory detoxification results. Due to the different decoding strategies employed by LMA compared to the other methods, direct comparisons of fluency and diversity are not entirely fair. Therefore, we only measured its detoxification metrics. The unsatisfactory performance can be attributed to the considerable parameter scale of LLMs. The Self-Detoxify method, which is tailored for smaller models such as GPT2-large, struggles with adaptation to LLMs. Moreover, the LMA method’s failure is due to the complex nature of toxicity across different LLM families. However, without finetuning or auxiliary models, our method still demonstrates competitive detoxification effects across these categories of LLMs, and our method performs well across different model families, proving its effectiveness and scalability.
Base Model Method Toxicity $\downarrow$ Fluency $\downarrow$ Diversity $\uparrow$
EMT TP PPL Dist-1 Dist-2 Dist-3
GPT2-XL Base 0.560 0.590 18.142 0.582 0.850 0.847
(1.3B) DeStein 0.322 0.160 24.989 0.592 0.865 0.859
Base 0.539 0.550 17.687 0.612 0.851 0.828
LLaMA2-7B Self-Detoxify 0.413 0.318 83.972 0.648 0.876 0.839
@@
DeStein 0.296 0.170 29.160 0.618 0.858 0.835
Base 0.622 0.661 16.127 0.565 0.839 0.841
OPT-6.7B Self-Detoxify 0.559 0.554 75.019 0.582 0.864 0.856
@@
DeStein 0.291 0.157 17.733 0.562 0.850 0.855
LLaMA2-13B Base 0.543 0.560 17.018 0.606 0.845 0.826
DAPT (LoRA) 0.473 0.440 20.424 0.593 0.836 0.824
DeStein 0.353 0.190 20.252 0.611 0.855 0.835
Table 3: Evaluation results on the RTP dataset with LLMs. The best results are shown in bold.
@@
LLM-as-a-Judge. We incorporate GPT-4 and Gemini (Durmus et al., 2022) as supplementary evaluators. We compare our method’s generative outputs against top-performing baselines: DisCup and DEXPERTS for GPT2-large, and Self-Detoxify and LMA for LLMs. For design details about the prompt, please see Appendix C. Based on our findings that the number of prompts did not significantly affect the results, we have extracted a balanced test dataset with 50 toxic and 50 non-toxic prompts for evaluation. The evaluation results indicate that our method exhibits lower toxicity and improved fluency compared to the best-performing baselines. The results are presented in Table 4.
Base Model Versus GPT-4 Gemini
Win Tie Lose Win Tie Lose
GPT2-Large DisCup 0.72 0.00 0.28 0.64 0.15 0.21
DEXPERTS 0.79 0.00 0.21 0.63 0.16 0.21
LLaMA2-7B Self-Detoxify 0.71 0.00 0.29 0.63 0.14 0.23
LMA 0.74 0.01 0.25 0.64 0.17 0.19
@@
Table 4: LLM-as-a-Judge evaluation results: Win, Lose, and Tie respectively represent the percentage of times our method outperforms, underperforms, or equals the baseline.
4.3 Ablation study
To analyze our method further, we study the contribution of different components on GPT2-large. Based on our findings that the number of prompts does not significantly affect the results, we have extracted a balanced test dataset consisting of 50 toxic and 50 non-toxic prompts for the ablation study. Our results, as shown in Table 5, demonstrate that removing self-induced data significantly raises toxicity (EMT: 0.327 to 0.203, TP: 0.19 to 0.04), whereas the removal of parallel data leads to a less pronounced increase in toxicity (EMT: 0.216, TP: 0.07). Regarding activation fusion, switching activation positions to FFN and FFN+MHSA notably increases PPL (148.14 and 59.848, respectively). These findings highlight the importance of self-induced parallel data and specific activation fusion techniques. Additionally, we observe a slight degradation in results when ’w/o head-wise coefficients’ is used. We offer a more detailed analysis in Appendix F.
Model Toxicity $\downarrow$ Fluency $\downarrow$ Diversity $\uparrow$
EMT TP PPL Dist-1 Dist-2 Dist-3
DeStein 0.203 0.04 39.405 0.569 0.858 0.860
Self-induced parallel pairs
w/o self-induced 0.327 0.19 32.145 0.566 0.862 0.863
w/o parallel 0.216 0.07 41.567 0.564 0.855 0.863
Activation fusion
w/o head-wise coefficients 0.207 0.04 39.434 0.569 0.859 0.860
Activation positions
FFN 0.404 0.26 148.14 0.576 0.867 0.868
FFN+MHSA 0.249 0.06 59.848 0.564 0.858 0.865
Table 5: Ablation study: w/o self-induced (using parallel pairs sourced from the ParaDetox dataset); w/o parallel (using non-parallel pairs from the self-induced dataset); w/o head-wise coefficients (using a one-size-fits-all fusion coefficient); activation positions (switching activation positions to FFN and FFN+MHSA). All results in the table are based on GPT2-large.
5 Further analysis
5.1 Trade-off between detoxification and task performance in LLMs
Existing methods of detoxification often focus on preserving linguistic quality in text generation. However, we emphasize extending evaluation criteria to broader language understanding and reasoning tasks essential for LLMs. Our approach incorporates MMLU (Massive Multitask Language Understanding) benchmark (Hendrycks et al., 2021) to evaluate the impact of detoxification techniques on task performance in LLMs. MMLU assesses how well a detoxified model retains its competence across multiple tasks, including question-answering, summarization, and sentiment analysis, providing a comprehensive measure of its practical utility. We show the evaluation results of task performance in Table 6.
The findings demonstrate that our method not only achieves exceptional detoxification effects but also maintains the task-solving abilities of LLMs. Specifically, our detoxification method and parameter-efficient finetuning method have a similar impact on the task-solving abilities of the LLaMA2-13b model.
Method Average weighed accuracy $\uparrow$ STEM $\uparrow$ Humanities $\uparrow$ Social sciences $\uparrow$ Other $\uparrow$
Random 0.250 0.250 0.250 0.250 0.250
Base 0.557 0.443 0.544 0.634 0.608
DAPT (LoRA) 0.530 0.437 0.493 0.612 0.592
DeStein 0.530 0.430 0.511 0.598 0.589
Table 6: Evaluation results of the trade-off between detoxification and task performance on the MMLU benchmark with LLaMA2-13b.
@@ 5.2 Analysis on the influence of detoxification strength
Figure 2: Trade-off between detoxification strength and PPL on GPT2-large.
For previous methods during the inference stage, when the detoxification strength ($\alpha_{contr}$) exceeds a certain critical value, it often leads to a rapid decrease in the fluency of generated text, rendering the text completely unusable. This limitation hinders the effectiveness of many methods in achieving efficient detoxification. However, our method detoxifies layer by layer through representations in activation spaces, which can largely avoid this problem. Our experiments, with results in Figure 2, demonstrate this.
We take the difference between the toxicity scores of the base model and the detoxified model as the detoxification strength and observe the relationship between fluency and detoxification strength. After reaching point A, the model’s generative capabilities only deteriorate significantly. However, by the time the control intensity reaches point A, the toxicity has already been reduced to 0.030. Such a low level of toxicity becomes imperceptible to humans. The range of control intensity before point A is already sufficient to meet the model’s detoxification needs, so the generative collapse beyond point A does not have an impact.
Previous decoding methods were largely based on the logit space, and as the control intensity increased, generation collapse became inevitable. The success of our experiment provides a new perspective for controllable generation that more efficient control may be achievable in activation spaces.
5.3 Analysis on interpretability in activation spaces
Controllable generation methods based on VAEs (Sohn et al., 2015) leverage a low-dimensional bottleneck to learn a disentangled latent space, while LMs often struggle to achieve good controllability at the activation layer due to the complexity of high-dimensional spaces. This proposed method mitigates the complexity of high-dimensional spaces to some extent through steering pairs and head-wise fusion. In this section, we verify the interpretability of the method by analyzing the data distribution in activation spaces.
@@ 5.3 Analysis on interpretability in activation spaces
As illustrated in Figure 3(a), we employ a heatmap to visualize the accuracy of linear classifiers for each attention head across all layers, and to analyze the effects and underlying principles of probing-based weights.
To further bolster the credibility of our method, we investigate the existence of a toxicity-non-toxicity vector within activation spaces. Specifically, we focus on the activation space corresponding to the 6th attention head of the 23rd layer, as illustrated in Figure 3(b). This layer and head were selected due to their high classification accuracy, which was identified through heatmap analysis. We employ PCA to visualize the representation of toxic and non-toxic samples within the activation space. Our results reaffirm the presence of a linear representation between toxicity and non-toxicity in the activation space. Moreover, to more vividly demonstrate the distribution differences in various activation spaces, we also visualize the space associated with the 12th layer and 7th head, as illustrated in Figure 3(c), which exhibits a classification accuracy nearly equivalent to that of random selection in the heatmap. The distribution of toxic and non-toxic samples within these two spaces further validates our method’s theoretical soundness.
@@ 6 Conclusions
In this paper, we propose a novel approach to toxicity mitigation, named DeStein, which is based on the activation space for detoxification. Specifically, our method utilizes the model’s self-induced data to identify the detoxification vectors. During the inference phase, the detoxification vectors and the output of the module are merged, and probe techniques are employed to apply weights to head-wise activations. Experimental results demonstrate that our approach not only achieves efficient detoxification but also preserves the generative capacity of the model to the greatest extent possible. Notably, DeStein maintains considerable performance in toxicity mitigation without significantly increasing inference time. Unlike previous methods, DeStein exhibits flexibility and competitiveness when dealing with LLMs and is scalable across different model families.
@@ 7 Limitaions
Our method, predicated on the assumption of linear representation, employs arithmetic operations to decouple toxic attributes and general capabilities, yet this approach is notably constrained. Initially, constructing an ideal parallel pair poses considerable difficulty. Moreover, this method represents an indirect form of decoupling, facing theoretical challenges in achieving complete separation. Consequently, there is a necessity to explore more efficient and versatile decoupling strategies, such as those based on causal reasoning, knowledge guidance, or meta-learning techniques. These methods aim to effectively unify general capabilities with safety, delineating our subsequent direction of research.